# DPMambaIR: All-in-One Image Restoration via Degradation-Aware Prompt State Space Model

### v2

@@ I Introduction
Image restoration is a fundamental task in computer vision, aiming to recover high-quality images from degraded inputs. Recent advances in deep learning have significantly improved performance in Degradation-specific restoration tasks, such as image denoising [1, 2], deblurring [3], deraining[4], dehazing[5, 6] and low-light enhancement[7]. However, traditional methods typically rely on Degradation-specific models explicitly tailored to individual degradation types. These approaches often require prior knowledge of the degradation type, which limits their practicality in real-world scenarios, such as autonomous driving and nighttime surveillance, where diverse and unknown degradations frequently coexist.
@@
To address this limitation, recent studies explore All-in-One image restoration frameworks that unify multiple restoration tasks into a single model. Existing methods can be broadly categorized into Mixture-of-Experts (MoE)-based and Prompt-based approaches. MoE-based methods, such as MoFME [8], dynamically route tasks to specialized experts using an uncertainty-aware router, enabling effective learning of diverse expert functions. However, these methods suffer from high computational costs, increased inference complexity as the number of experts grows, and lack explicit degradation modeling, limiting adaptability to unseen or complex degradations. Prompt-based methods, like OneRestore [9], guide restoration using degradation prompts, integrating scene descriptors via cross-attention mechanisms for flexible and adaptive strategies. While effective, these methods rely on coarse prompts that fail to capture nuanced degradation characteristics and treat prompts as static inputs, limiting their ability to adapt to varying degradation intensities or specific requirements, leading to suboptimal results.
While these methods have demonstrated success in learning multiple tasks for All-in-One image restoration, two key challenges remain unresolved. First, existing models lack a degradation extractor capable of capturing fine-grained degradation features and often rely on coarse degradation labels to guide the restoration process. For example, images with varying degrees of haze are typically assigned a single haze label, failing to capture the nuanced variations in degradation levels. This simplification leads to suboptimal restoration performance, as the model cannot adapt to the diverse characteristics of different degradations. Second, current methods struggle to balance the competing requirements of varying restoration tasks. For instance, training low-light enhancement, which focuses on restoring low-frequency contrast information, alongside deblurring, which prioritizes the recovery of fine local details. This inherent task competition makes it challenging to optimize a single network for all restoration tasks simultaneously and remains a fundamental challenge for All-in-One image restoration frameworks.
To address these challenges, we propose DPMambaIR, a novel All-in-One image restoration framework built upon a regression-based degradation extractor and a Degradation-Aware Prompt State Space Model (DP-SSM). Unlike prior approaches that use static or class-based prompts, our method employs a degradation extractor to generate prompt vectors that encode both semantic categories and degradation distribution characteristics. These prompts are then used to dynamically modulate the state transition, input, and output matrices of a state space model, enabling fine-grained degradation modeling and adaptive coordination across frequency bands. This mechanism allows our model to better manage task competition and enhance restoration quality across diverse degradations. Moreover, while State Space Model such as Mamba exhibit strong global modeling capabilities, they are inherently biased toward low-frequency restoration [10]. To better recover local high-frequency details, we introduce a lightweight High-Frequency Enhancement Block (HEB) as a supplementary component, which delivers performance gains while incurring only negligible computational overhead.
@@
Extensive experiments on a mixed dataset containing seven degradation types demonstrate the superiority of DPMambaIR. As shown in Fig. 1, DPMambaIR achieves state-of-the-art performance in terms of PSNR and SSIM, outperforming existing All-in-One methods such as AdaIR and OneRestore. Moreover, DPMambaIR achieves competitive results across individual tasks, including deraining, low-light enhancement, deblurring, and dehazing, showcasing its potential as an effective and robust All-in-One image restoration solution.
@@
    We propose a novel degradation extractor capable of capturing fine-grained degradation features from complex degraded images. These features provide critical priors to improve the performance of All-in-One image restoration.
@@
    We design a Degradation-aware Prompt State Space Model (DP-SSM), which enhances the long-range modeling capability of Mamba by integrating degradation priors at multiple structural levels.
@@
    Extensive experiments on a mixed dataset containing seven types of degradation demonstrate that DPMambaIR achieves state-of-the-art performance on PSNR and SSIM metrics, outperforming existing methods. The results validate the effectiveness and superiority of the proposed approach.
@@ II-A All-in-One Image Restoration
Image restoration is a fundamental task in computer vision, aiming to recover high-quality images from degraded inputs. Traditional methods, such as Dark Channel Prior [11] and Color Line Prior [12], rely on handcrafted priors to model degradation but struggle with complex scenarios due to predefined assumptions. The rise of deep learning [13] has significantly advanced tasks like such as image super-resolution[14, 15, 16], image denoising [1, 2], deblurring [3], deraining[4], dehazing[5], desnowing[17] and low-light enhancement[7]. These approaches, leveraging convolutional neural networks (CNN)-based [18], transformer-based [19], or diffusion-based [20] architectures, enable powerful feature learning for image restoration. However, most existing works focus on addressing a single type of degradation, which limits their applicability in real-world scenarios where images often suffer from diverse and unknown types of degradation.
Recent advancements in All-in-One image restoration frameworks focus on enhancing flexibility and efficiency by leveraging various multi-task learning techniques such as Mixture-of-Experts (MoE) and text-based or vision-based prompts to handle diverse restoration tasks. For example, AirNet[21] employs contrastive learning to extract degradation representations for guiding restoration. PromptIR[22] integrates end-to-end prompt generation and interaction to adaptively handle different degradation types, while OneRestore[9] aligns text-based degradation categories with visual features as prompts. IDR[23] uses meta-learning to decompose degradations, enabling generalization across diverse scenarios. MoFME[8], on the other hand, employs an uncertainty-aware router to dynamically assign Degradation-specific features to expert modules, allowing it to specialize in multiple tasks. Similarly, AdaIR[24] achieves adaptive restoration by decoupling degradations in the frequency domain. Additionally, methods like DA-CLIP[25] leverage the pre-trained vision-language model (VLM) for prompt extraction, but the significant computational overhead introduced by VLM limits their practical efficiency.
Despite their effectiveness, most prompt-based methods rely on coarse-grained class labels or static degradation tokens, which limits their ability to model continuous degradation distributions or adapt to image-specific degradation characteristics. In contrast, our method extracts fine-grained, regression-level degradation embeddings that encode both semantic and statistical degradation features. These embeddings are used to dynamically modulate the state-space parameters of the restoration backbone, enabling degradation-aware processing throughout the network. This facilitates better coordination across restoration tasks and reduces competition between frequency components.
@@ II-B State Space Model based Image Restoration
In Image Restoration, Convolutional Neural Networks (CNN) [13] and Transformer [19] have been widely adopted due to their strong feature extraction and representation learning capabilities. CNN excels at capturing local patterns, making them suitable for tasks requiring high spatial resolution and fine details. However, their limited receptive field often hinders their ability to model long-range dependencies, which are critical for handling global degradations. On the other hand, Transformer leverages self-attention mechanisms to capture global context effectively. Despite their success, Transformers suffer from quadratic complexity relative to the input size, making them computationally and memory-intensive.
To address these limitations, researchers have turned to the State Space Model (SSM)[26], which provides an efficient framework for modeling long-range dependencies with linear computational complexity. Formally, SSM maps an input sequence $x(t)$ to an output sequence $y(t)$ through an implicit latent state $h(t)\in\mathbb{R}^{N}$, represented by a linear ordinary differential equation:
@@
where $A\in\mathbb{R}^{N\times N}$, $B\in\mathbb{R}^{N\times 1}$, $C\in\mathbb{R}^{1\times N}$, and $D\in\mathbb{R}$. After discretization with step size $\Delta$, the dynamic equation can be expressed as:
@@
where $\overline{A}=\exp(\Delta A)$ and $\overline{B}=(\Delta A)^{-1}(\exp(\Delta A)-I)\Delta B$.
The introduction of Vision Mamba[27] marked the first effort to incorporate SSM into vision tasks. By proposing the Vision State Space Module (VSSM), it extended SSM to visual applications and demonstrated superior performance over Vision Transformer (ViT)[28], while maintaining lower complexity. This success inspired extensive research into SSM across diverse tasks such as object detection[29], semantic segmentation[30], and image classification[31].
In the context of Image Restoration, MambaIR[32] pioneered the integration of SSM, leveraging its linear complexity and long-range modeling capability to achieve significant improvements in both computational efficiency and perceptual quality. This was followed by several variants targeting specific restoration tasks[33, 34, 35], collectively showcasing the strong potential of SSM in this domain. Existing approaches typically unfold 2D images into 1D sequences and apply sequential causal modeling. However, this design still faces two limitations: (a) Limited Global Context Utilization: The $i$-th pixel can only access the previous i1i-1 pixels in the sequence, restricting the effective use of global context. Parallel multi-directional scanning has been explored to enlarge the receptive field, but it increases computational cost while providing limited gains in low-level vision tasks[32]. (b) Lack of Effective Degradation-Aware Mechanisms: Current SSM-based methods primarily target single-task restoration. Their transition and output mappings are fixed or solely input-driven, without leveraging external degradation priors. As a result, they cannot dynamically adapt to diverse degradation types or levels, leading to suboptimal performance in All-in-One image restoration task.
These limitations motivate the introduction of a degradation-aware modulation mechanism into the SSM pipeline. In this work, we propose a Degradation-aware Prompt State Space Model (DP-SSM), where degradation embeddings dynamically modulate the state-space transition and output matrices. This design enables a fine-grained, task-aware mechanism that balances global modeling with degradation-specific guidance, enhancing the flexibility and effectiveness of SSM for complex restoration scenarios.
@@ III-A Problem Definition
$\boldsymbol{I}_{D}=Deg(\boldsymbol{I}_{C})+\boldsymbol{\eta}.$ (4)
Here, $\boldsymbol{I}_{C}\in\mathbb{R}^{3\times H\times W}$ denotes the original clean image, $Deg$ represents the degradation operator (e.g., blur kernel, brightness, or quality compression), and $\boldsymbol{\eta}$ signifies additive noise. The main challenge arises from the fact that $Deg$ is usually unknown and varies across tasks, making it difficult to design a universal model capable of effectively handling all possible degradations.
@@ III-B DPMambaIR
We propose DPMambaIR, a degradation-aware framework for All-in-One image restoration. As illustrated in Fig. 2, we first establish a Mamba-based image restoration baseline using an asymmetric encoder-decoder design within a U-shaped architecture. The restoration process begins with a $3\times 3$ convolutional layer, which extracts shallow features $\boldsymbol{F}\in\mathbb{R}^{H\times W\times C}$ from the degraded image. These features then pass through three encoding stages, each followed by downsampling. During each encoding stage, the resolution of the feature map is halved while the number of channels is doubled. Subsequently, the features pass through three decoding stages, where each stage involves upsampling and concatenating the output of the decoder with the corresponding output of the encoder to form the input for the next decoding stage. Finally, after a refinement stage, the clean image is restored through a shortcut connection that combines the input degraded image with the output of the decoder.
To address the challenges posed by unknown degradation types and levels in All-in-One image restoration tasks, we propose a Degradation Extractor pre-trained using degraded image reconstruction. This model compresses degradation type and severity information into a compact embedding, enabling the model to transition from blind to non-blind image restoration through finer-grained degradation representation. To further optimize the utilization of the extracted degradation embedding, we revisit the selective state-space modeling process. Specifically, we propose a Degradation-Aware Prompt State Space Model, which guides the image restoration process by compressing historical information and supplementing global degradation information.
@@ III-C Degradation Extractor
To better extract degradation information from images, we revisit the degradation modeling process of real-world images, which spans from image capture and imaging to transmission, as illustrated in Fig. 3 (a). This process typically involves four common types of degradations: (1) poor illumination conditions, (2) relative motion between the target objects and the camera, (3) artifacts introduced by transmission media in the optical path, and (4) compression during storage and transportation. We unify these scenarios into a generalized degradation formation model:
@@
where $J(x)$ represents the clean image, $\alpha(x)$ models the luminance adjustment, such as the effects of insufficient lighting in low-light or adverse weather conditions (e.g., rain or fog), and $\gamma(x)$ accounts for artifacts caused by occlusions from aerosols, raindrops, or snow particles in the transmission medium. $\beta(x)$ modulates $\gamma(x)$ based on varying spatial properties, $K$ represents the blur introduced by relative motion between the target and the camera, and $Q$ denotes the compression applied during storage and transmission.
@@
Based on the aforementioned degradation modeling process, we design a universal degradation extractor, as illustrated in Fig. 3 (b). Specifically, we introduce central difference convolution (CDC)[36, 37, 38] to extract edge and texture degradation cues related to illumination and blur. In addition, we propose a multi-scale degradation perception module, which employs four parallel CDC operators of different sizes to capture degradation features at various scales. By concatenating the outputs, the module retains complete degradation representations across all frequency components. Furthermore, we employ three groups of CDC with different scales to capture artifacts of varying sizes, such as visual artifacts and snowflakes. Additionally, we integrate a traditional convolution branch to learn brightness-related representations.
To achieve a balance between extraction performance and computational cost, we introduce a re-parameterization technique for efficient integration of degradation representations, as illustrated in Fig. 3 (c). These two implementations correspond to different variants of our method, which are presented in Table I as DPMambaIR-L and DPMambaIR, respectively.
Furthermore, to endow the degradation extractor with fine-grained degradation encoding capability, we propose a pretraining strategy based on degraded image reconstruction. This strategy leverages a hybrid dataset composed of paired samples with diverse degradation types. The key objective is to train the degradation extractor $\mathcal{E}$ to capture detailed degradation information from a degraded image $\boldsymbol{I}_{D}$ and compress it into a compact embedding $\boldsymbol{E}_{d}$, formulated as:
$\boldsymbol{E}_{d}=\mathcal{E}(\boldsymbol{I}_{D}).$ (6)
Subsequently, the reconstruction unit $\mathcal{R}$ takes $\boldsymbol{E}_{d}$ together with the clean content embedding $\boldsymbol{E}_{c}$ to reconstruct the degraded image $\boldsymbol{I}_{D}$. In particular, we adopt Restormer as the content extractor to derive $\boldsymbol{E}_{c}$ from the clean reference image $\boldsymbol{O}$, expressed as:
$\boldsymbol{E}_{c}=Restormer(\boldsymbol{O}),\ \ \boldsymbol{\hat{I}}_{D}=\mathcal{R}(\boldsymbol{E}_{c},\boldsymbol{E}_{d}).$ (7)
This pretraining paradigm encourages the degradation extractor to encode both the type and severity of degradation into a continuous embedding space, which then serves as a compact and informative task descriptor for downstream restoration. We further introduce a dual-component loss function $L$, which combines the $L_{1}$ norm and the Learned Perceptual Image Patch Similarity (LPIPS) metric. The combined loss function is expressed as $L(\boldsymbol{I}_{D},\boldsymbol{\hat{I}}_{D})=L_{1}+L_{LPIPS}$, striking a balance between reconstruction accuracy and perceptual quality.

### v3

@@ I Introduction
Image restoration is a fundamental task in computer vision, aiming to recover high-quality images from degraded inputs. Recent advances in deep learning have significantly improved performance in degradation-specific restoration tasks, such as image denoising [64, 19], deblurring [66], deraining[38], dehazing[43, 47], and low-light enhancement[30]. However, traditional methods typically rely on degradation-specific models explicitly tailored to individual degradation types. These approaches often require prior knowledge of the degradation type, which limits their practicality in real-world scenarios, such as autonomous driving and nighttime surveillance, where diverse and unknown degradations frequently coexist.
@@
To address this limitation, recent studies explore All-in-One image restoration frameworks that unify multiple restoration tasks into a single model. Existing methods can be broadly categorized into Mixture-of-Experts (MoE)-based and Prompt-based approaches. MoE-based methods, such as MoFME [63], dynamically route tasks to specialized experts using an uncertainty-aware router. However, these methods suffer from high computational costs and lack explicit degradation modeling, limiting adaptability to unseen degradations. Alternatively, Prompt-based methods like OneRestore [11] guide restoration using degradation prompts, integrating scene descriptors via cross-attention mechanisms.
However, existing prompt-based strategies [11, 39] suffer from fundamental limitations in how degradation information is represented and exploited. A key shortcoming lies in their reliance on static priors, where degradations are encoded as discrete class labels or fixed semantic tokens. Such categorical representations are inherently inadequate for modeling the continuous nature of real-world degradations and their spatial variability, for example, differentiating between light haze and dense, spatially accumulating fog. Furthermore, most of these methods follow an additive fusion paradigm in which degradation cues are treated merely as auxiliary information appended to visual features. This mechanism is intrinsically limited, as the underlying network architecture remains governed by fixed convolutional kernels or static attention patterns. Consequently, the model is compelled to address heterogeneous restoration demands, such as suppressing high-frequency noise while simultaneously enhancing global contrast, within a rigid geometric structure. This conflict ultimately degrades performance across diverse and complex degradation scenarios.
To address these challenges, we propose DPMambaIR, a novel All-in-One framework that shifts the design philosophy from conventional feature fusion to explicit dynamic parameter modulation. Our approach is built upon a fine-grained degradation extractor and a Degradation-Aware Prompt State Space Model (DP-SSM). Unlike methods that rely on discrete degradation labels, the proposed extractor employs a reconstruction objective to regress continuous degradation embeddings, enabling it to encode nuanced information regarding both degradation type and severity. These embeddings are utilized to directly modulate the core State Space Model parameters $(\Delta,B,C)$. Modulating the discretization step $\Delta$ effectively alters the integration step size of the underlying ordinary differential equation, allowing the model to realize distinct continuous dynamics conditioned on degradation characteristics. Specifically, the network learns to assign smaller $\Delta$ values to induce a high-inertia state favorable for smoothing high-frequency noise, whereas larger values are generated to enter a high-gain regime that amplifies weak signals in low-light conditions. Concurrently, the input projection matrix $B$ is modulated to control how strongly external observations drive the latent state transitions, while the output matrix $C$ is adjusted to inject global degradation priors. Furthermore, we observe that jointly optimizing heterogeneous degradations often induces an implicit bias toward low-frequency structures, making the consistent recovery of fine details challenging. To alleviate this issue, we incorporate a lightweight High-frequency Enhancement Block (HEB) that complements the proposed dynamic modulation and facilitates the restoration of local textures.
@@
Extensive experiments on a mixed dataset containing seven degradation types demonstrate the efficacy of DPMambaIR. As shown in Fig. 1, DPMambaIR achieves state-of-the-art performance in terms of PSNR and SSIM, outperforming existing All-in-One methods such as AdaIR and OneRestore. Moreover, DPMambaIR achieves competitive results across individual tasks, including deraining, low-light enhancement, deblurring, and dehazing, demonstrating its potential as an effective and robust All-in-One image restoration solution.
@@
    We propose a degradation extractor capable of capturing fine-grained, continuous degradation features from complex degraded images via a regression-based reconstruction objective.
@@
    We design a Degradation-aware Prompt State Space Model (DP-SSM) that introduces a dynamic parameter modulation mechanism. By dynamically modulating the state-space evolution parameters based on degradation priors, it enhances the model’s physical adaptability to diverse degradation dynamics.
@@
    Extensive experiments on a mixed dataset containing seven types of degradation demonstrate that DPMambaIR achieves state-of-the-art performance on PSNR and SSIM metrics, outperforming existing methods. The results validate the effectiveness and robustness of the proposed approach.
@@ II-A All-in-One Image Restoration
Image restoration remains a foundational challenge in computer vision, striving to reconstruct high-fidelity content from corrupted inputs. While traditional approaches, such as Dark Channel Prior [13] and Color Line Prior [7], relied on handcrafted heuristics that often struggle in complex scenarios, the advent of deep learning [21] has revolutionized the field. Significant strides have been made in specific domains, including super-resolution [24, 36, 37], denoising [64, 19], deblurring [66, 61, 62], deraining [38, 59], dehazing [43, 17], desnowing [10, 60], and low-light enhancement [30]. Leveraging CNN-based [22], Transformer-based [45], or diffusion-based [14] architectures, these methods excel at feature representation. However, their reliance on degradation-specific designs fundamentally limits their utility in real-world environments where degradations are diverse, compounded, and often unknown.
Recent research has consequently pivoted towards All-in-One frameworks that emphasize flexibility and unified processing. These methods generally employ multi-task learning strategies, such as Mixture-of-Experts (MoE) or prompt-guided adaptation. For instance, AirNet [23] utilizes contrastive learning to derive degradation representations, while MoFME [63] implements an uncertainty-aware router to dynamically dispatch features to specialized experts. Although effective, MoE architectures often incur high computational overhead and lack explicit physical modeling of the degradation. In parallel, prompt-based methods such as PromptIR [39], OneRestore [11], and DA-CLIP [27] integrate degradation cues, ranging from learned visual prompts to text embeddings, via cross-attention mechanisms. IDR [58] further explores meta-learning for decomposition, and AdaIR [4] adopts frequency domain decoupling for adaptive restoration.
Despite these advancements, fundamental limitations persist in the representation and exploitation of degradation information. Most prompt-based strategies [39, 11] rely on coarse-grained priors or discrete class labels, which fail to capture the continuous nature and spatial variability inherent in real-world images. Furthermore, they typically adopt an additive fusion paradigm where degradation cues are treated merely as auxiliary inputs. This mechanism is intrinsically limited because the underlying network parameters remain fixed, governed by static weights or attention patterns that lack the flexibility to adapt to heterogeneous restoration demands. In contrast, our DPMambaIR framework employs a regression-based extractor to derive fine-grained, continuous degradation embeddings. We depart from simple feature concatenation and utilize these embeddings to dynamically re-parameterize the State Space Model. This shift from additive fusion to explicit dynamic parameter modulation allows the system to adapt its evolution dynamics to the specific physical properties of the degradation.
@@ II-B State Space Model based Image Restoration
Convolutional Neural Networks (CNNs) [21] and Transformers [45] have long dominated image restoration. While CNNs excel at capturing local high-frequency details, their limited receptive fields hinder the modeling of long-range dependencies essential for correcting global degradations like haze or uneven lighting. Transformers address this with global self-attention but suffer from quadratic computational complexity relative to image resolution.
To reconcile long-range modeling with efficiency, researchers have revisited the State Space Model (SSM) [18]. SSMs map a 1D input sequence $x(t)$ to an output $y(t)$ via a latent state $h(t)\in\mathbb{R}^{N}$, governed by the linear ordinary differential equation (ODE):
@@
where $A,B,C,D$ are learnable parameters. For digital implementation, this continuous system is discretized with a step size $\Delta$, yielding the recurrence:
@@
Here, the discretized matrices are defined as $\overline{A}=\exp(\Delta A)$ and $\overline{B}=(\Delta A)^{-1}(\exp(\Delta A)-I)\Delta B$. This formulation allows for linear-complexity inference while maintaining a global receptive field.
The Vision Mamba [65] first adapted SSMs for computer vision, demonstrating that the Vision State Space Module could outperform Vision Transformers [6] with significantly lower computational costs. This success has sparked adoption across tasks including object detection [50], segmentation [49], and classification [42]. In image restoration, MambaIR [9] pioneered the use of SSMs to achieve a balance between efficiency and perceptual quality, inspiring subsequent variants [8, 35, 5].
However, applying standard SSMs to All-in-One restoration faces two primary limitations. First, Limited Global Context Utilization: The causal nature of standard scanning means the $i$-th pixel only accesses information from preceding pixels. While multi-directional scanning can alleviate this, it increases computation without always yielding proportional gains in low-level tasks [9]. Second, and more critically, Lack of Degradation-Awareness: Existing SSM-based restoration methods rely on fixed transition dynamics ($A,B,C$) or purely input-driven modulation. They lack an explicit mechanism to adapt the system’s evolution rules based on external degradation priors (e.g., adjusting the integration step $\Delta$ for different noise levels). Consequently, their ability to generalize across the diverse degradation types encountered in All-in-One settings remains constrained.
Motivated by these gaps, we propose the Degradation-aware Prompt State Space Model (DP-SSM). Unlike prior works, our method utilizes fine-grained degradation embeddings to dynamically modulate the discretization step $\Delta$, along with the state-space transition and output matrices. This design establishes a dynamics-aware mechanism that harmonizes global context modeling with degradation-specific physical guidance, significantly enhancing robustness in complex restoration scenarios.
@@ III-A Problem Definition
$\boldsymbol{I}_{D}=Deg(\boldsymbol{O})+\boldsymbol{\eta}.$ (4)
Here, $\boldsymbol{O}\in\mathbb{R}^{3\times H\times W}$ denotes the original clean image, $Deg$ represents the degradation operator (e.g., blur kernel, brightness, or quality compression), and $\boldsymbol{\eta}$ signifies additive noise. The main challenge arises from the fact that $Deg$ is usually unknown and varies across tasks, making it difficult to design a universal model capable of effectively handling all possible degradations.
@@ III-B DPMambaIR
We propose DPMambaIR, a degradation-aware framework for All-in-One image restoration. As illustrated in Fig. 2, our method is built upon a Mamba-based U-shaped architecture with an asymmetric encoder-decoder design. Initially, a $3\times 3$ convolution extracts shallow features $\boldsymbol{F}\in\mathbb{R}^{H\times W\times C}$ from the degraded input. The encoder comprises three stages, each utilizing downsampling to progressively halve the feature resolution while doubling the channel dimension. Correspondingly, the decoder employs upsampling and skip connections to fuse multi-scale features from the encoding path. Finally, the clean image is reconstructed via a refinement stage, utilizing a global residual connection to superimpose the learned residual onto the degraded input.
To address the challenges posed by unknown degradation types and severities in All-in-One image restoration, we propose a Degradation Extractor pre-trained via a self-supervised reconstruction objective. This module encodes heterogeneous degradation patterns into a compact embedding, effectively bridging the gap between blind and non-blind restoration by providing explicit degradation representations. To fully exploit these representations, we revisit the selective state-space modeling paradigm. Specifically, we propose a Degradation-Aware Prompt State Space Model, which dynamically modulates the restoration trajectory by injecting global degradation priors into the state transition process.
@@ III-C Degradation Extractor
To effectively extract degradation priors, we formulate a generalized degradation model that encapsulates the physical corruption process from image capture to transmission, as shown in Fig. 3 (a). This process typically involves four degradation categories: (1) illumination degradation, (2) motion-induced blur, (3) transmission medium artifacts, and (4) compression artifacts. We unify these factors into a composite formation model:
@@
where $J(x)$ denotes the latent clean image. The term $\alpha(x)$ models luminance variations, encompassing low-light conditions or uneven illumination caused by weather. $\gamma(x)$ represents additive artifacts introduced by the transmission medium, such as rain streaks, snow, or haze aerosols, while $\beta(x)$ modulates the spatial intensity of these artifacts. $K$ denotes the blur kernel resulting from relative motion, and $Q$ represents quantization or compression operations applied during storage.
Guided by this formulation, we design a degradation extractor capable of capturing heterogeneous corruption patterns, as illustrated in Fig. 3(b). Specifically, we employ Central Difference Convolution (CDC) [46, 33, 34] to identify gradient-based degradation cues, such as edges and textures associated with blur and noise. A multi-scale module aggregates features from four parallel CDC branches with varying kernel sizes, enabling the perception of degradations across different frequency bands. Additionally, we integrate a standard convolution branch to capture global intensity and brightness distributions.
To balance extraction accuracy with computational efficiency, we employ a re-parameterization technique during inference, merging the multi-branch structure into a compact representation, as shown in Fig. 3 (c). This yields two model variants: a larger capacity DPMambaIR-L utilizing the full multi-branch structure, and a lightweight DPMambaIR employing the re-parameterized extractor (detailed in Table I).
To enable the learning of fine-grained degradation embeddings, we propose a reconstruction-based pre-training strategy. Unlike classification-based methods that rely on discrete labels, this approach forces the extractor to learn a continuous manifold of degradation types and severities. Formally, the extractor $\mathcal{E}$ maps a degraded input $\boldsymbol{I}_{D}$ to a low-dimensional embedding $\boldsymbol{E}_{d}$:
$\boldsymbol{E}_{d}=\mathcal{E}(\boldsymbol{I}_{D}).$ (6)
To supervise this process, a reconstruction decoder $\mathcal{R}$ attempts to recover the degraded image $\boldsymbol{I}_{D}$ by combining $\boldsymbol{E}_{d}$ with a content embedding $\boldsymbol{E}_{c}$ derived from the corresponding clean ground truth $\boldsymbol{O}$:
$\boldsymbol{E}_{c}=\text{Restormer}(\boldsymbol{O}),\ \ \boldsymbol{\hat{I}}_{D}=\mathcal{R}(\boldsymbol{E}_{c},\boldsymbol{E}_{d}).$ (7)
This disentanglement ensures that $\boldsymbol{E}_{d}$ captures pure degradation information necessary to reconstruct the corruption, independent of the image content. The optimization minimizes a hybrid objective combining $L_{1}$ loss and the Learned Perceptual Image Patch Similarity (LPIPS) metric, defined as $\mathcal{L}=\|\boldsymbol{I}_{D}-\boldsymbol{\hat{I}}_{D}\|_{1}+\lambda\mathcal{L}_{\text{LPIPS}}$, to ensure perceptually faithful reconstruction of the degradation patterns.